Armidale University offers only high-tech graduate-level programs. Armidale has two principal operating departments, Engineering and Computer Sciences, and two support departments, Facility and Technology Maintenance and Enrollment Services. The base used to allocate Facility and Technology Maintenance is budgeted total maintenance hours. The base used to allocate Enrollment Services is number of credit hours for a department. The Facility and Technology Maintenance budget is $500,000, while the Enrollment Services budget is $1,200,000. The following chart summarises budgeted amounts and allocation- base amounts used by each department:

Services Provided (Annually)

Budget

Engineering

Computer Sciences

F&T Maintenanc e

Enrollment Services

F&T Maintenance(in hours)

$500,000

1,000

2,000

Zero

5,000

Enrollment Services
(in credit hrs)

$1,200,000

24,000

36,000

2,000

Zero

Required:
Prepare a schedule which allocates service department costs using the step-down method with the sequence of allocation based on the highest-percentage support concept. Compute the total amount of support costs allocated to each of the two principal operating departments, Engineering and Computer Sciences.
Answer #1

In companies

  • the operating departments are the main ones- eg production department in manufacturing companies
  • Support function is done by service departments

Service department costing is done to allocate the cost of service departments to operating departments.

There are several methods for this

  • Direct cost of a service department is allocated to operating departments
  • step down etc

cost of a service department is allocated to other service departments as well as operating departments

STEP DOWN

Cost of a service department is allocated to other service departments as well as operating departments.

The allocation is done in a sequential manner

Here there are

  • two service department-F&T Maintenance and Enrollment Services
  • two operating department- Engineering and Computer Sciences

Total service provided by each service departments

  • F&T Maintenance(in hours)   -           8,000
  • Enrollment Services (in credit hrs)    -         62,000

Percentage support provided

Among service provided by service departments to other service departments, which one ranks the highest in terms of percentage ?

Support provided by F&T to enrollment- 5000 hrs

Percentage support by F&T to enrollment= 5000/8000 =0.6250

(62.5% of F&T's service is provided to enrollment)

Support provided by enrollment to F&T - 5000 hrs

Percentage support by enrollment to F&T = 2000/62000 =0.0323

(3.23% of enrollment's service is provided to F&T )

The service provided by F&T to enrollment is greater than The service provided by enrollment to F&T

F&T provides the largest amount of its service to other support departments. So the cost of F&T is allocated first.

F&T Maintenance $500,000

to Enrollment Services = $500,000 x 5000/8000=3,12,500 $

to Engineering = $500,000 x 1000/8000= $ 62,500

to Computer Science = $500,000 x 2000/8000= $ 1,25,000

Now total costs of Enrollment Service = $12, 00,000 + 3,12,500 = $15,12,500

This cost is allocated to Engineering and Computer Science

to Engineering = $15,12,500 x 24000/60000 = $667500

to Computer Science = $15,12,500 x 36000/60000 = $81032500

Sevice departments

Operating departments

Before allocation

F&T Maintenanc e

Enrollment Services

Engineering

Computer Sciences

Department costs

500000

1200000

Allocation

from F& T others

-500000

312500

62500

125000

from enrolment to others

-1512500

605000

907500

Total

Total costs after alloxation

0

0

667500

1032500

1700000